A small table of values is presented below for sel
Answer #1

1)

We use Snell's law of refraction

n1 sin theta 1 = n2 sin theta 2

n2 = n1 * sin theta 1/ sin theta 2

n2 = sin 30 / sin 15 = 1.93

2)

We use Snell's law of refraction

n1 sin theta 1 = n2 sin theta 2

n2 = n1 * sin theta 1/ sin theta 2

n2 = 1.33 *sin 45 / sin 30 = 1.88

3)

Since it moves from a denser to rarer medium, Speed INCREASES

4)

C

Decreases to keep Frequency constant, as velocity decreases

5)

B

Frequency Remains the Same